NASA International Space Station On-Orbit Status 25 January 2017

©NASA

NASA International Space Station On-Orbit Status 25 January 2017.

Today: Fluid Shifts Chibis Imaging Day 2: The Flight Day 45 Chibis session for the 49S crew was planned this morning, however due to Joint Station LAN (JSL) network issues the session had to be cancelled. This Fluid Shifts session will be replanned for tomorrow.

Lighting Effects Vision Test: The crew performed a Numerical Verification and a Color Discrimination test in their Crew Quarters with illumination from the Solid State Lighting Assembly (SSLA). The Lighting Effects experiment hopes to better quantify and qualify how lighting can effect habitability of spacecraft. The light bulbs on the ISS are being replaced with a new system designed for improved crew health and wellness. The Lighting Effects investigation studies the impact of the change from fluorescent light bulbs to solid-state light-emitting diodes (LEDs) with adjustable intensity and color and aims to determine if the new lights can improve crew circadian rhythms, sleep, and cognitive performance. Results from this investigation also have major implications for people on Earth who use electric lights.

Habitability Software Update: The crew installed updated software on their tablets in support of the Habitability experiment. The iQ&A iPad application will be used to collect questionnaire data for Habitability Experiment. The Habitability investigation collects observations about the relationship between crew members and their environment on the ISS. Observations can help spacecraft designers understand how much habitable volume is required, and whether a mission's duration impacts how much space crew members need.

Internal Thermal Control System (ITCS) Sample Collection: The crew conducted ITCS fluid sampling from the Node 2 (N2) Moderate Temperature Loop (MTL) and Low Temperature Loop (LTL) sample ports. They also sampled Columbus ITCS water using the Columbus Coolant Fluid Servicer. These samples will be returned to the ground on SpX-10 for analysis. The return to ground samples are taken once a year to monitor the quality of the water in the ITCS loops.

Major Constituent Analyzer (MCA) Mass Spectrometer Assembly Remove and Replace (R&R): On GMT 2016/286, the Node 3 MCA experienced a shutdown due to a failure in its Mass Spectrometer Assembly. A spare Mass Spectrometer Assembly was launched on HTV-6, and it was used today to replace the failed MCA Mass Spectrometer Assembly in Node 3 (N3).

Mobile Servicing System (MSS) Cupola Artificial Vision Unit (AVU) Troubleshooting: Crew set up a video camcorder such that Cupola Remote Workstation (RWS) Mon 1 and 2 are in the video frame. Ground specialists routed the camcorder and AVU video output to the ground, the Cupola AVU was then commanded from the ground to power-up perform a checkout. Several different inputs to the AVU (test pattern, external cameras, both SCU sources) were tried during the checkout, and all resulted in mostly blank displays with small portions of the display occasionally flickering into view. It was concluded that the input to the AVU is not the issue. The problem is likely in the AVU itself. This checkout was to determine the viability of utilizing the computing capacity of this MSS equipment last used for ISS operations in 2007.

H-II Transfer Vehicle 6 (HTV6) Departure Preparation: Yesterday the crew completed loading more than 1450 kg of trash into the HTV vehicle. Today they performed a review of the HTV6 departure profile, procedures, and crew interfaces for monitoring and commanding HTV6. They also performed Robotics training which consisted of HTV6 release scenarios to allow them to practice techniques for nominal and off nominal situations. Additionally, they removed General Luminaire Assemblies (GLAs) from the forward side of HTV, and Light Emitting Diode (LED) Lights from the aft side of HTV in preparation for HTV Departure. These will all be stowed on ISS and used as spares. HTV6 is scheduled to unberth from the ISS on Friday 27-January.

Today's Planned Activities
All activities were completed unless otherwise noted.

Fine Motor Skills Experiment Test - Subject
FLUID SHIFTS. Comm configuration on RS
A/C Inverter Swap in SM
FLUID SHIFTS. Camcorder setup & switch on
Crew Prep for PAO Event
Ultrasound 2 HRF Rack 1 Power On
Total Organic Carbon Analyzer (TOCA) Water Recovery System (WRS) Sample Analysis
Fluid Shifts Ultrasound 2 Data Export Alternate
TV-session with the participants of the 11th All-Russia Youth Readings named after S.P. Korolev
Habitability iPad App Update [Deferred]
Fluid Shifts OCT Service Module Setup Alternate
Atmosphere Control and Supply (ACS) Nitrogen Manual Valve Close
FLUID SHIFTS. Подключение Laptop OCT к БРИ и HRF РС к ППС-26 (вместо RSE-Med) [Deferred]
Node 2 Low Temperature Loop (LTL) Internal Thermal Control System (ITCS) Sample Collection
СОЖ maintenance
Fluid Shifts OCT Service Module Power On
Fluid Shifts CCFP Service Module Configuration
Fluid Shifts DPOAE Service Module Setup
Node 2 Moderate Temperature Loop (MTL) Internal Thermal Control System (ITCS) Sample Collection
FLUID SHIFTS. Don Chibis [Deferred]
FLUID SHIFTS. Assistance with Chibis & Gamma-1 [Deferred]
Fluid Shifts Tonometry Service Module Setup
Fluid Shifts Experiment Service Module Data Collection [Deferred]
Lab Low Temperature Loop (LTL) Internal Thermal Control System (ITCS) Sample Collection
FLUID SHIFTS. Doff Chibis [Deferred]
FLUID SHIFTS. Disconnection & disassembly [Deferred]
Fluid Shifts CCFP DPOAE Service Module End [Deferred]
FLUID SHIFTS. Comm Reconfiguration to nominal on RS
Fluid Shifts Tonometry Service Module Stow
Delta file prep
Fluid Shifts OCT Service Module Power Off
FLUID SHIFTS. Отключение Laptop OCT, HRF и подключение RSE-Med [Deferred]
Total Organic Carbon Analyzer (TOCA) Sample Data Record
FLUID SHIFTS. DROUTE [Deferred]
Fluid Shifts OCT Service Module Stow
Fluid Shifts CCFP HRF PC Service Module Stow
Fluid Shifts Hardware Final Gather [Deferred]
Node 3 Major Constituent Analyzer (MCA) Mass Spectrometer (MS) Assembly Remove and Replace
Fluid Shifts Hardware USOS Transfer [Deferred]
Actiwatch Spectrum HRF1 Configuration Alternate 2
Life On The Station Photo and Video
WIRELESS LEAK DETECTOR ULTRASONIC SENSOR REMOVAL, FILE TRANSFER, AND STOW
Health Maintenance System (HMS) OCT Setup
Taking Pille dosimeter readings
Robotic Workstation (RWS) High Definition (HD) Monitor Downlink
Stow CWSA samples, units and tools.
Robotic Workstation (RWS) Setup
Lighting Effects Visual Performance Tests Light Setting - Subject
Restoring COL1D1 configuration
Marrow Breath And Ambient Air Sample Setup - Subject
Columbus ITCS water sample using Columbus Coolant Fluid Servicer for Return to Ground
WIRELESS LEAK DETECTOR ULTRASONIC SENSOR
Robotic Work Station (RWS) Artificial Vision Unit (AVU) Troubleshooting
Health Maintenance System (HMS) - OCT Exam
Health Maintenance System (HMS) OCT Stow
Remote Workstation Monitor HD Video Downlink Deactivation RWS.
Internal Thermal Control System (ITCS) Sample Stow
Major Constituent Analyzer (MCA) - Hand Valve 01 (HV01) Open
Health Maintenance System (HMS) Fundoscope Setup
Major Constituent Analyzer (MCA) - Hand Valve 01 (HV01) Close
Health Maintenance System (HMS) Fundoscope Prep
ESA Active Dosimeter Mobile Unit Swap
Health Maintenance System (HMS) Eye Exam - Fundoscope
Health Maintenance System (HMS) Fundoscope Tear Down and Stow

Completed Task List Items
MD Print Removal and Stow
Columbus 16-Port Ethernet Switch Audit

Ground Activities
All activities were completed unless otherwise noted.
RWS AVU Checkout
Node 3 CDRA Activation
Node 3 MCA Activation

Three-Day Look Ahead:
Thursday, 01/26: HTV5 (Depart/ROBOT OBTs, Hatch Closure, Vestibule Demate, CPA Install), Eye Exams, ARED Detent R&R
Friday, 01/27: HTV Release and Departure, SPHERES Zero Robotics Run
Saturday, 01/28: Weekly Housekeeping, Crew Off Duty

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - Off
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - On
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Operate
Major Constituent Analyzer (MCA) Node 3 - Idle
Oxygen Generation Assembly (OGA) - Standby
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Full Up
Trace Contaminant Control System (TCCS) Node 3 - Off

Please follow SpaceRef on Twitter and Like us on Facebook.